# Interpretable factorization of clinical questionnaires to identify latent factors of psychopathology

### 12 Dec 2023 v1

@@ 5 Discussion
In this paper, we introduced ICQF, a non-negative matrix factorization method designed for questionnaire data. Our method incorporates characteristics that enhance the interpretability of the resulting factorization, as conveyed by psychiatry collaborators. We showed that their qualitative desiderata can be turned into formal constraints in the factorization problem, together with direct modelling of confounding variables, which other methods do not allow. The method is user friendly, by supporting automated estimation of the number of factors, minimizing the number of hyper-parameters, and transparently handling missing entries instead of requiring separate imputation. The characteristics above mean that ICQF required an entire optimization procedure to be derived from scratch. We provided a theoretical formalization of the problem and the procedure, and demonstrated a pair of propositions that guarantee convergence of the procedure to a local minimum and, in certain conditions, a global minimum as well.

### 17 Mar 2026 v2

@@ 5 Discussion
In this paper, we introduced ICQF, a non-negative matrix factorization method designed for questionnaire data. Our method incorporates characteristics that enhance the interpretability of the resulting factorization, as conveyed by psychiatrist collaborators. We showed that their qualitative desiderata can be turned into formal constraints in the factorization problem, together with direct modelling of confounding variables, which other methods do not allow. The method is user friendly, by supporting automated estimation of the number of factors, minimizing the number of hyper-parameters, and transparently handling missing entries instead of requiring separate imputation. The characteristics above mean that ICQF required an entire optimization procedure to be derived from scratch. We provided a theoretical formalization of the problem and the procedure, and demonstrated a pair of propositions that guarantee convergence of the procedure to a local minimum and, in certain conditions, a global minimum as well.